OpenStudy (anonymous):

Find a function that describes the sequence. Use y for the term in the sequence and n for its place in the list. Then use the function to find the next three terms in the sequence. 21, 22, 23, 24, . . . a.y = 20n +1; 101, 121, 141 c.y = 21n; 105, 126, 147 b.y = n + 20; 25, 26, 27 d.y = 22n –1; 109, 131, 153

OpenStudy (anonymous):

b. y=n+20; 25,26,27 because if you plug in 1 into the equation it equals 21. 2+20=22, etc. so to find the next three terms of the equation plug in 5: y=5+20=25; plug in 6: y=6+20=26; and plug in 7: y=7+20=27

OpenStudy (anonymous):

Find a function that describes each sequence. Use y for the term in the sequence and n for its place in the list. Then use the function to find the next 3 terms in the sequence. 3, 6, 9, 12,. . . a. y = 6n - 3; 27, 33, 39 c. y = n + 3; 15, 18, 21 b. y = 3n; 15, 18, 21 d. y = 3n + 3; 18, 21, 24

OpenStudy (anonymous):

b. y=3n; 15,18,21. Plug 1 in: 3(1)=3; 3(2)=6;etc. to find the next 3 terms: plug in 5: y=3(5)=15; plug in 6: y=3(6)=18; plug in 7: y=3(7)=21